OpenStudy (anonymous):

write the first five terms for the sequence an=n(n+1)

OpenStudy (anonymous):

find value of an for n=1, n=2, n=3, n=4 and n=5

OpenStudy (anonymous):

is it 2,5,7,9,11

OpenStudy (ghazi):

nope

OpenStudy (ghazi):

check out your second term

OpenStudy (ghazi):

and all the remaining terms

OpenStudy (ghazi):

putting n=2 gives you an= 2*(2+1)

OpenStudy (anonymous):

4

OpenStudy (ghazi):

nope ..it's 2*3=6

OpenStudy (anonymous):

how

OpenStudy (anonymous):

nvm lol

OpenStudy (ghazi):

solve the brackets first then multiply the outcome of bracket by2 like 2+1=3 and multiplied by 2=6

OpenStudy (anonymous):

i get it idk what i was looking at

OpenStudy (anonymous):

a1=2 a2=2*3=6 a3=3*4=12 a4=4*5=20 a5=5*6=30

OpenStudy (anonymous):

what about this an=(-2)^n

OpenStudy (anonymous):

what do I do

OpenStudy (ghazi):

just put n=1,2,3,4,5 and so on.....like (-2)^1=-2 but in case of even power result will be positive

OpenStudy (anonymous):

so sombody told me its -2,4,-8,16,-32 but i dont understand when you plug those numbers in you dont get those unless you plug in the answer of what you got before

OpenStudy (anonymous):

a1=(-2)^1=-2 a2=(-2)^2=4 a3=(-2)^3=8 . . . . .

OpenStudy (amistre64):

make sure you have the proper starting value of "n"; some texts say 1, others say 0

OpenStudy (anonymous):

omg idk what im thinking im so lost today lol thanks really

OpenStudy (anonymous):

ohhh sry..... a3=(-2)^3=-8

OpenStudy (anonymous):

Yah i got it i was thinking times each other not to the power

OpenStudy (anonymous):

Now for something like this what do I do same question as befoe a2=10, d=-2